# TEASMA: A Practical Approach for Test Assessment of Deep Neural Networks using Mutation Analysis

### v2

@@ II-D Mutants and Faults
In traditional software, MOs are defined to introduce simple syntactical changes representing real mistakes often made by programmers into the source code of a program. Similarly, DeepCrime defined a set of pre-training MOs based on an analysis of real faults [17, 18, 19] related to the training set or program used to train a DNN model. However, the high cost of generating mutants using pre-training MOs hinders their application in practice. Post-training MOs, in contrast, apply modifications directly to the original model but have been criticized for performing modifications that are not realistic [22]. Nevertheless, for mutation analysis to be effective in assessing the adequacy of test sets, regardless of how mutation operators are defined, it is important to generate mutants whose corresponding MS is, for a test set, a good predictor of fault detection. Therefore, considering the strong practical advantages of post-training MOs, despite their presumed lack of realism, they should be carefully investigated to build predictors of fault detection before being dismissed as inadequate.
@@
The use of mutation analysis for assessing the quality of test sets in traditional software is supported by the observed association between mutants and real faults, as reported in the literature [15, 16, 8]. This association has been investigated by performing correlation analysisbetween the MS of a test set and its capability to detect faults. Regression analysis has also been applied for modeling this relationship [15, 8]. In DNNs, however, the association between mutants and faults has never been investigated. The main challenge in such investigation is that, in contrast to traditional software where it is possible to identify real faults by isolating the faulty statement that caused a failure, identifying faults in DNNs is not a straightforward task due to the complexity and non-linear nature of the DNNs. However, a recent approach proposed by Aghababaeyan et al. [6] for estimating faults is a possible, practical, and automated solution. We describe next how we rely on this approach to investigate the relationship between mutants and faults, and how MS can help predict FDR.
II-E Fault Estimation
@@
The number of mispredicted test inputs is a well-known metric in evaluating DNN testing approaches [1, 23, 24]. However, the number of mispredictions is not a good indicator of the number of faults and hence it is not an adequate metric to perform mutation analysis. Indeed, a single fault within a DNN model can make the model mispredict multiple test inputs. Subsequently, a test set that identifies a considerable number of mispredictions may actually detect only a few underlying faults, whereas another test set with the same number of mispredictions could detect a higher number of distinct faults. Therefore, assessing test sets based on mispredictions and thus investigating the relationship between the number of killed mutants and mispredictions can be misleading.
@@
Consequently, we rely on a recent approach introduced by Aghababaeyan et al. [6] for estimating faults in a DNN model, that are defined as distinct root causes of mispredictions [25]. Aghababaeyan et al. [6] cluster mispredicted inputs that exhibit similar features in three key steps: feature extraction, dimensionality reduction, and density-based clustering. Initially, VGG-16 [26] is employed to extract the feature matrix from the mispredicted inputs, serving as the foundation for clustering. Subsequently, dimensionality reduction is applied to enhance clustering performance within the inherently high dimensional feature space. The HDBSCAN algorithm [27] is then utilized to cluster the mispredicted inputs based on the extracted features. Aghababaeyan et al. [6] empirically analyzed the resulting clusters and showed that inputs in the same cluster are mispredicted due to the same fault in the model and, conversely, inputs in different clusters are mispredicted due to distinct faults. Therefore, if a test set contains only one of the test inputs in a cluster, then it is able to detect the underlying fault of that cluster.
@@
III The TEASMA Approach
In this section, we present TEASMA, a practical solution for assessing test sets for DNNs using mutation analysis. The basic idea is to rely on mutation and regression analysis based on the training set to accurately predict the FDR of test sets, which can then be used for evaluating their adequacy before proceeding with labeling and execution. We provide an overview of the main steps to apply TEASMA in Figure 2 and a detailed description of these steps in Algorithms1 and 2. The process to be followed by engineers in practice is illustrated in Figure 2 and includes three high-level steps:
1) Calculating MS and FDR for subsets of the training set.
In this step, only post-training MOs are used to generate mutants as described in Section II-A, since in our application context pre-training MOs are not a practical option. Faults of the DNN model are estimated using the misprediction clustering approach described in Section II-E based on the training input set. Then, a large number of subsets of varying sizes are sampled from the training set, and their MS and FDR are calculated to be used in the next step.
@@ III The TEASMA Approach
In this step, regression analysis between FDR and MS is performed to obtain the best regression model for predicting FDR based on MS. This prediction model is then used in the next step to assess the adequacy of test sets.
@@
In this step, given a test set, its MS is calculated using the original DNN model and previously generated mutants. Then, the prediction model obtained from the previous step is used to predict the test set’s FDR ($\widehat{FDR}$) along with Prediction Intervals (PI).
The Algorithm 1 describes the first two steps implemented in TEASMA to create an accurate regression model for predicting FDR based on MS. As illustrated in Figure 2, the first two steps of TEASMA are applied entirely on the training set of the DNN. The algorithm starts by initializing an empty archive of subset samples $A$ (line 1). Then, to be able to perform mutation analysis, TEASMA is required to generate mutants (line 2) and identify faults in the DNN (line 3). TEASMA leverages DeepMutation++ to generate and select the set of mutants $M$ (Section III-A). It also relies on a SoTA approach to identify the set of distinct faults $F$ in the DNN model based on the model’s mispredicted inputs in the training set (Section III-B).
Input: $N$: the original DNN model under test $TrSet$: the labeled training input set of $N$ $SN$: number of samples per each sampling size
Output: $PM$: prediction model built for N
$sample\ archive\ A\xleftarrow{}\emptyset$   Generate post-training mutants $mutants\ M\xleftarrow{}GenerateMutants(N);$   Estimate faults based on the training set $faults\ F\xleftarrow{}EstimateFautls(N,TrSet);$   1) Calculate MS and FDR for training input subsets do
1       $size\xleftarrow{}UpdateSamplingSize(A);$ $samples\xleftarrow{}Sampling(TrSet,size,SN,type);$ foreach subset $sub_{i}$ in $samples$ do
2             $MS_{i}\xleftarrow{}MSCalc(sub_{i},M,N,MSFormula);$ $FDR_{i}\xleftarrow{}FDRCalc(sub_{i},F);$
3      $A\xleftarrow{}AddtoArchive(samples,MS,FDR);$
while $(0\leq FDR\leq 1\ is\ not\ covered)$  2) Perform regression analysis and select best model $regression\ models\ RM\xleftarrow{}\emptyset$ $RM\xleftarrow{}RegressionAnalysis(A);$ $PM\xleftarrow{}argmax(RM,R^{2});$ return $PM$
@@ 3      $A\xleftarrow{}AddtoArchive(samples,MS,FDR);$
The algorithm includes a main input sampling loop for selecting a large number of input subsets with various sizes from the training set of the original DNN model $TrSet$ and for calculating their MS and FDR (lines 4-11). The sampling is performed until the entire FDR range ($0\leq FDR\leq 1$) is covered. In each iteration, TEASMA updates the sampling size (line 5) with a larger or smaller size according to the minimum and maximum FDR achieved by the samples currently present in the archive $A$. At the first iteration, when the archive is empty, the $UpdateSamplingSize$ function starts with a small $size$. To be able to build a regression model, TEASMA requires multiple input subsets of the same size. Therefore, for each subset $size$, TEASMA samples $SN$ subsets from the training set (line 6). $SN$ is one of the input parameters of the algorithm and can be defined based on the size of the training set. For the reasons we describe later in Section III-C, TEASMA can sample input subsets either randomly or uniformly. After sampling input subsets, TEASMA calculates the MS and FDR of all subsets (lines 7-9). The MS of each subset can be computed based on one of the MS calculation approaches ($MSFormula$) described in Section II-B. In the following sections, we investigate the relationship between FDR and MS, as calculated using each MS calculation approach, to determine the best approach for building an accurate FDR prediction model. FDR is then computed based on the formula specified in Section II-E. Finally, the MS and FDR corresponding to all sampled subsets are added to the archive $A$ (line 10).
@@
With a large number of input subsets of diverse sizes selected from the training set and their corresponding MS and FDR, TEASMA builds a set of regression models (lines 12 and 13). For reasons we will explain in Section III-D, TEASMA considers non-linear regression models including quadratic, exponential, and regression trees in addition to linear regression. To enable the engineer to identify the best regression model, TEASMA relies on common metrics for evaluating such models (Section III-D) including the coefficient of determination ($R^{2}$) to measure the goodness of fit. Using these metrics, engineers can determine for each DNN and training input set, the best regression model for predicting FDR based on MS. As a default, TEASMA selects the regression model with the highest goodness of fit ($R^{2}$) (line 14).
Input: $T$: an unlabeled test set to be assessed $N$: the original DNN model under test
Output: $\widehat{FDR}(T)$: predicted FDR of $T$ PI: prediction interval of $\widehat{FDR}(T)$
  Load mutants generated for N $M\xleftarrow{}LoadMutants(N)$   Load the prediction model built for N $PM\xleftarrow{}LoadPredictionModel(N)$   3) Assess test set T based on its MS $MS_{T}\xleftarrow{}MSCalculation(T,M,N,MSFormula);$ $[\widehat{FDR},PI]\xleftarrow{}Predict(PM,MS_{T})$ return [$\widehat{FDR}\ ,\ PI$]

### v3

@@ II-A4 Mutants and Faults
In traditional software, MOs are defined to introduce simple syntactical changes representing real mistakes often made by programmers into the source code of a program. Similarly, DeepCrime defined a set of pre-training MOs based on an analysis of real faults [22, 23, 24] related to the training set or program used to train a DNN model. However, the high cost of generating mutants using pre-training MOs hinders their application in practice. Post-training MOs, in contrast, apply modifications directly to the original model but have been criticized for performing modifications that are not realistic [15]. However, it is important to note that not all faults in DNN models stem from human error [2]. As emphasized by Ma et al. [2], post-training operators are not intended to directly replicate human faults but rather to provide quantitative measurements of test set quality. Nevertheless, for mutation analysis to be effective in assessing the adequacy of test sets, regardless of how mutation operators are defined, it is important to generate mutants whose corresponding MS is, for a test set, a good predictor of fault detection. Therefore, considering the strong practical advantages of post-training MOs, we decided to use these operators in our experiments. Although these operators are not derived from the analysis of human faults, we investigate their potential as predictors of fault detection and thus as a tool to assess the adequacy of test sets.
@@
The use of mutation analysis for assessing the quality of test sets in traditional software is supported by the observed association between mutants and real faults, as reported in the literature [16, 17, 12]. This association has been investigated by performing correlation analysis between the MS of a test set and its capability to detect faults. Regression analysis has also been applied for modeling this relationship [16, 12]. In DNNs, however, the association between mutants and faults has never been investigated. The main challenge in such investigation is that, in contrast to traditional software where it is possible to identify real faults by isolating the faulty statement that caused a failure, identifying faults in DNNs is not a straightforward task due to the complexity and non-linear nature of the DNNs. However, a recent approach proposed by Aghababaeyan et al. [18] for estimating faults is a possible, practical, and automated solution. We describe in Section II-D how we rely on this approach to investigate the relationship between mutants and faults, and how MS can help predict FDR.
II-B Surprise Coverage
Surprise Coverage (SC) metrics have been introduced by Kim et al. [7, 6] as innovative test adequacy metrics that quantify the surprise or novelty of test inputs with respect to the training data of a DNN. These metrics are derived from the Surprise Adequacy (SA) concept, which evaluates how surprising a test input is to a DNN based on its training set, thereby indicating the potential for revealing unexpected behavior or faults. SC evaluates the diversity of a test set by measuring the distribution of SA among its test inputs relative to the training input set. Kim et al. [6] define three SA metrics for each test input:
Distance-based SA (DSA) utilizes the Euclidean distance between the activation traces of test inputs and those observed during training. This metric is particularly useful in classification tasks, where inputs closer to decision boundaries are deemed more valuable for testing due to their higher potential to detect misclassifications.
Likelihood-based SA (LSA) leverages Kernel Density Estimation (KDE) [25] to calculate the likelihood of the activation patterns of the test inputs given the distribution of training data activation patterns. Lower likelihoods suggest higher surprise, implying that the DNN is less familiar with such inputs, which might lead to higher fault revelation.
Mahalanobis Distance SA (MDSA) employs the Mahalanobis distance to measure the difference between activation traces of the test input and the distribution of the activation traces derived from the training inputs. This method is beneficial due to its sensitivity to both the mean and covariance of the training data distribution, providing a robust indication of input novelty.
Kim et al. [6] define three SC metrics: Likelihood-based Surprise Coverage (LSC), Distance-based Surprise Coverage (DSC), and Mahalanobis Distance Surprise Coverage (MDSC), which rely on LSA, DSA, and MDSA metrics, respectively. However, they only implemented LSC and DSC in their public repository [6]. The SC computation involves dividing the observed SA values range into distinct intervals or buckets. SC is then defined as the proportion of these buckets containing at least one test input, measuring the test set’s ability to cover a wide range of model behaviors:
$SC(T)=\frac{|\{b_{i}\mid\exists t\in T:SA(t)\in b_{i}\}|}{n}$ (5)
where $T$ is the set of test inputs, $b_{i}$ represents discrete buckets of SA values, and $n$ is the total number of buckets. In our experiments, we rely on the original implementation of the LSC and DSC metrics provided by Kim et al. [6].
II-C Input Distribution Coverage
Input Distribution Coverage (IDC), introduced by Dola et al. [5], is a black-box coverage metric that employs a novel approach by utilizing Variational Autoencoders (VAEs) to model the input distribution of a DNN, offering a framework to assess test adequacy based on the diversity and distribution of features within the test set. For each input set, a VAE is trained on its training input set, which is representative of the operational domain of any DNN trained with this input set. The IDC framework utilizes disentangled latent representations to learn a latent space composed of linear subspaces, where variation in each subspace corresponds to variation in a feature of the input set. However, methods for identifying the features present in a latent representation and accurately mapping these features to latent subspaces are not available. Consequently, the number of features representing an input space, i.e., the latent space dimensionality, is defined by a parameter in IDC.
The VAE is trained to encode the high-dimensional input space into a lower-dimensional latent space, enabling the IDC framework to define a feasible coverage domain. Consequently, the latent space is divided into distinct partitions, meaning any latent linear subspace is encompassed by a subset of partitions. These partitions represent unique regions that could potentially exhibit different behaviors or characteristics relevant to the DNN model’s performance. While testing the full combinatorial space is not practical, IDC relies on Combinatorial Interaction Testing (CIT) to measure test coverage. IDC thus assesses the adequacy of a test set based on how thoroughly its test inputs cover the diverse partitions within the latent space.
When assessing a test set, the encoder part of the VAE is used to obtain the latent representation of its test inputs. However, utilizing a VAE’s latent representation is only possible for inputs that follow the same distribution as the VAE’s training set, i.e., in-distribution inputs. Consequently, the IDC framework utilizes SOTA out-of-distribution (OOD) detectors to filter out OOD test inputs before processing them with the VAE. As a result, unlike all white-box DNN coverage metrics that measure test coverage for both in-distribution and OOD test inputs, IDC measures coverage solely based on in-distribution inputs. White-box DNN coverage metrics such as neuron coverage and extended neuron coverage cannot differentiate between valid and invalid test inputs and thus suffer from falsely inflated coverage by OOD inputs [26]. IDC overcomes the inclusion of OOD test inputs and enables a more accurate calculation of test coverage.
II-D Fault Estimation
@@
The number of mispredicted test inputs is a well-known metric in evaluating DNN testing approaches [8, 7, 27]. However, the number of mispredictions is not a good indicator of the number of faults and hence it is not an adequate metric to perform mutation analysis [18]. Indeed, a single fault within a DNN model can make the model mispredict multiple test inputs. Subsequently, a test set that identifies a considerable number of mispredictions may actually detect only a few underlying faults, whereas another test set with the same number of mispredictions could detect a higher number of distinct faults. Therefore, assessing test sets based on mispredictions and thus investigating the relationship between the number of killed mutants and mispredictions can be misleading.
@@
Consequently, we rely on a recent approach introduced by Aghababaeyan et al. [18] for estimating faults in a DNN model, that are defined as distinct root causes of mispredictions [28]. Aghababaeyan et al. [18] cluster mispredicted inputs that exhibit similar features in three key steps: feature extraction, dimensionality reduction, and density-based clustering. Initially, VGG-16 [29] is employed to extract the feature matrix from the mispredicted inputs, serving as the foundation for clustering. Subsequently, dimensionality reduction is applied to enhance clustering performance within the inherently high dimensional feature space. The HDBSCAN algorithm [30] is then utilized to cluster the mispredicted inputs based on the extracted features.
Aghababaeyan et al. [18] used both manual and metric-based evaluation to select the best hyperparameter configuration from several clustering configurations. For the metric-based evaluation, they used two established metrics: Silhouette score [31] and Density-Based Clustering Validation (DBCV) [32]. To complement these quantitative assessments, they conducted a manual analysis to assess the quality of the final selected clusters. For this purpose, they generated heatmaps for each cluster, visualizing the distribution of features across inputs within a cluster. Their empirical analysis of the heatmaps showed that inputs in the same cluster are mispredicted due to the same fault in the model and, conversely, inputs in different clusters are mispredicted due to distinct faults. Therefore, if a test set contains only one of the test inputs in a cluster, then it is able to detect the underlying fault of that cluster.
Similar failure clustering approaches have been proposed by Biagiola and Tonella [33] and Attaoui et al. [34] with different objectives. Biagiola and Tonella [33] used the clustering approach to evaluate the diversity of the generated inputs while Attaoui et al. [34] relied on the clusters for test selection aiming to efficiently improve the performance of the DNN model through retraining. Unlike Aghababaeyan et al. [18], given their objectives, neither of these approaches proposed a fault analysis procedure to evaluate whether clusters capture faults. Consequently, in this study, we rely on the clustering approach introduced by Aghababaeyan et al. [18] and adopt the same manual and metric-based approach to evaluate the identified fault clusters.
@@
Figure 2: The process of test adequacy assessment with TEASMA
III The TEASMA Methodology
In this section, we present TEASMA, a methodology for assessing the adequacy of test sets by predicting their fault detection capability, based on the training set and utilizing existing adequacy metrics. The basic idea is to rely on regression analysis and the training set to accurately predict the FDR of test sets, which can then be used for evaluating their adequacy before proceeding with labeling and execution. TEASMA provides the flexibility to perform adequacy assessment using any metric but one needs to determine which one yields the most accurate FDR predictions. In this study, we apply TEASMA with four SOTA test adequacy metrics (MS, DSC, LSC, and IDC), which were described in Section II. We then investigate later in Section V which of these metrics can predict FDR more accurately. We provide an overview of the main steps to apply TEASMA in Figure 2 and a detailed description of these steps in Algorithms 1 and 2. The process to be followed by engineers in practice is illustrated in Figure 2 and includes three high-level steps:
1) Calculating Adequacy Score (AS) and FDR for subsets of the training set.
In this step, faults of the DNN model are estimated using the misprediction clustering approach described in Section II-D based on the training input set. Then, a large number of subsets of varying sizes are sampled from the training set, and their AS and FDR are calculated to be used in the next step. To calculate the AS of these training subsets, the selected adequacy metric needs to be configured. The configuration depends on the selected adequacy metric. As mentioned earlier, we repeat TEASMA using MS, SC, and IDC for calculating AS in this study. Therefore, AS configuration in our experiments includes mutant generation, configuring SC parameters, and training a VAE using the training set. The AS configuration for the selected metric is saved to be reused by the last step of TEASMA.
@@ III The TEASMA Methodology
In this step, regression analysis between FDR and AS is performed to obtain the best regression model for predicting FDR based on AS. This prediction model is then used in the next step to assess the adequacy of test sets. In this study, we investigate four different adequacy metrics, and thus perform regression analysis separately for FDR and each metric. The optimal regression model is then determined in terms of FDR prediction accuracy.
@@
In this step, given a test set, its AS is calculated based on the selected adequacy metric and the same configuration as in previous steps. For instance, MS is calculated using the original DNN model and previously generated mutants. Similarly, SC is computed using the same configuration, including the DNN layer selected for computing SA, the number of buckets, and the lower and upper bounds of SA values, as in the first step. For IDC, the same pre-trained VAE is used. Then, using the computed AS, the prediction model obtained from the previous step is used to predict the test set’s FDR ($\widehat{FDR}$) along with Prediction Intervals (PI).
The Algorithm 1 describes the first two steps implemented in TEASMA to create an accurate regression model for predicting FDR. As illustrated in Figure 2, the first two steps of TEASMA are applied entirely on the training set of the DNN. The algorithm starts by initializing an empty archive of subset samples $A$ (line 1). Then, to be able to perform regression analysis, TEASMA is required to configure the AS computation (line 2) and identify faults in the DNN (line 3). The configuration process depends on the selected adequacy metric. In this study, it includes either generating mutants for computing MS, configuring the parameters for SC calculation, or training a VAE for IDC calculation.
Input:
$N$: the original DNN model under test
$TrSet$: the labeled training input set of $N$
$SN$: number of samples per each sampling size
1
Output:
$PM$: prediction model built for N
2
[…]
4
  Configure AS calculation
5 $config\xleftarrow{}AdequacyConfig();$
6
  Estimate faults based on the training set
[…]
8
  1) Calculate MS and FDR for training input subsets
9 do
10       $size\xleftarrow{}UpdateSamplingSize(A);$
11       $samples\xleftarrow{}Sampling(TrSet,size,SN);$
12       foreach subset $sub_{i}$ in $samples$ do
13             $AS_{i}\xleftarrow{}AdequacyCalc(sub_{i},N,config)$
14             $FDR_{i}\xleftarrow{}FDRCalc(sub_{i},F);$
15
[…]
17
18while $(1-maxFDR\geq\theta)\>and\>(minFDR\geq\theta)$
  2) Perform regression analysis and select best model
19 $regression\ models\ RM\xleftarrow{}\emptyset$
20 $RM\xleftarrow{}RegressionAnalysis(A);$
21 $PM\xleftarrow{}argmax(RM,R^{2});$
22 return $PM$
23
@@ 21 $PM\xleftarrow{}argmax(RM,R^{2});$
The algorithm includes a main input sampling loop that selects a large number of input subsets with various sizes from the training set of the original DNN model $TrSet$, calculates their AS and FDR (lines 4-11), and adds them to the archive $A$. At the first iteration, when the archive is empty, the $UpdateSamplingSize$ function starts with a relatively small $size$. In each iteration, TEASMA updates the sampling size (line 5) with a larger or smaller size according to the minimum and maximum FDR achieved by the samples currently present in the archive $A$. The sampling is performed until the entire FDR range ($0\leq FDR\leq 1$) is covered. To achieve this, we define a stopping criterion based on the minimum ($minFDR$) and maximum ($maxFDR$) FDR values achieved by the samples in $A$ and a threshold parameter $\theta$. If $1-maxFDR\geq\theta$, the $UpdateSamplingSize$ function increases the sampling size to enable sampling subsets that are slightly larger than the largest subset in $A$ and can achieve higher FDR. Similarly, this function decreases the sampling size to enable sampling subsets slightly smaller than the smallest subset in $A$, if $minFDR\geq\theta$. This process is repeated with various subset sizes until $(1-maxFDR<\theta)\>and\>(minFDR<\theta)$. We set $\theta=0.05$ for our experiments in this study.
To be able to build a regression model, TEASMA requires multiple input subsets of the same size. Therefore, for each subset $size$, TEASMA samples $SN$ subsets from the training set (line 6). $SN$ is one of the input parameters of the algorithm and can be defined based on the size of the training set. For the reasons we describe later in Section III-C, input subsets can be sampled either randomly or uniformly. After sampling input subsets, TEASMA calculates the AS based on the selected metric and the FDR of all subsets (lines 7-9). When using MS, it can be computed based on any of the MS calculation approaches described in Section II-A3. In the following sections, we investigate the relationship between FDR and each AS to determine the best metric for building an accurate FDR prediction model. FDR is then computed based on the formula specified in Section II-D. Finally, the AS and FDR corresponding to all sampled subsets are added to the archive $A$ (line 10).
@@
With a large number of input subsets of diverse sizes selected from the training set and their corresponding AS and FDR, TEASMA builds a set of regression models (lines 12 and 13). For reasons we will explain in Section III-D, we consider non-linear regression models including quadratic, exponential, and regression trees in addition to linear regression. To identify the best regression model, TEASMA relies on common metrics for evaluating such models (Section III-D), including the coefficient of determination ($R^{2}$) to measure the goodness of fit. Using these metrics, engineers can determine, for each DNN and training input set, the best regression model for predicting FDR. As a default, TEASMA selects the regression model with the highest goodness of fit ($R^{2}$) (line 14).
Input:
$T$: an unlabeled test set to be assessed
$N$: the original DNN model under test
1
Output:
$\widehat{FDR}(T)$: predicted FDR of $T$
PI: prediction interval of $\widehat{FDR}(T)$
2
  Load AS calculation configuration for N
3 $config\xleftarrow{}LoadAdequacyConfig();$
4
  Load the prediction model built for N
5 $PM\xleftarrow{}LoadPredictionModel(N)$
6
  3) Assess test set T based on its AS
7 ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{AS_{T}\xleftarrow{}AdequacyCalc(T,N,config);}}$
8 $[\widehat{FDR},PI]\xleftarrow{}Predict(PM,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{AS_{T}}})$
9
10return [$\widehat{FDR}\ ,\ PI$]